# “I am bad”: Interpreting Stealthy, Universal and Robust Audio Jailbreaks in Audio-Language Models

### v1

@@ Abstract
The rise of multimodal large language models has introduced innovative human-machine interaction paradigms but also significant challenges in machine learning safety. Audio-language Models (ALMs) are especially relevant due to the intuitive nature of spoken communication, yet little is known about their failure modes. This paper explores audio jailbreaks targeting ALMs, focusing on their ability to bypass alignment mechanisms. We construct adversarial perturbations that generalize across prompts, tasks, and even base audio samples, demonstrating the first universal jailbreaks in the audio modality, and show that these remain effective in simulated real-world conditions. Beyond demonstrating attack feasibility, we analyze how ALMs interpret these audio adversarial examples and reveal them to encode imperceptible first-person toxic speech - suggesting that the most effective perturbations for eliciting toxic outputs specifically embed linguistic features within the audio signal. These results have important implications for understanding the interactions between different modalities in multimodal models, and offer actionable insights for enhancing defenses against adversarial audio attacks.
@@ 1 Introduction
Large Language Models (LLMs) have proven useful beyond a doubt across various domains since their widespread deployment, significantly enhancing productivity in tasks such as natural language processing, code generation, and creative content creation (Brown, 2020; OpenAI, 2024). However, their vast capabilities pose a considerable challenge in balancing usefulness and harmlessness (Bommasani, 2022). One prominent aspect of AI safety is alignment: ensuring that generated content corresponds to the functional objectives and ethical ideals of human users, minimizing risks of harm, bias, or misuse in real-world applications (Weidinger et al., 2021; Russell, 2022). Despite the development of various methods for alignment, such as reinforcement learning from human feedback (Christiano et al., 2023) and rule-based constraints (Mu et al., 2024), LLM alignment has been shown to be inherently brittle and easy to circumvent using adversarial prompts, jailbreak techniques, or context manipulation (Perez et al., 2022; Liu et al., 2024; Wei et al., 2023; Xu et al., 2024).
Humans naturally interact with their surroundings not only through written word, but more commonly via visual and spoken cues. This motivates the development of multimodal models, which integrate information from various modalities - such as text, images, and audio - to more effectively simulate human-like understanding and improve interaction with users (Baltrušaitis et al., 2019). A plethora of multimodal models have been released in recent years (Alayrac et al., 2022; Liu et al., 2023; Driess et al., 2023). Two such kinds of models are Vision Language Models (VLMs) (Zhang et al., 2024) and Audio Language Models (ALMs) (Chu et al., 2023), which respectively take image or audio and textual input simultaneously. Naturally, the introduction of an additional input channel opens a conspicuous attack vector through which a model can be deceived into undesirable output (Eykholt et al., 2018a; Jia et al., 2022). Adversarial example generation for plain language models is computationally expensive due to the discrete nature of the input space - specifically, gradient-based optimization methods do not directly map onto valid textual tokens, making it difficult to manipulate the input effectively (Carlini & Wagner, 2018). In contrast, when working with continuous signals, gradient manipulations can directly influence the input, enabling the generation of adversarial examples without the constraints imposed by discrete token boundaries.
As Vision-Language Models (VLMs) have become mainstream, with numerous commercial and open-source implementations available such as BLIP (Li et al., 2022), Flamingo (Alayrac et al., 2022) and CLIP (Radford et al., 2021), there has been extensive research into various types of attacks targeting the visual modality (Goodfellow et al., 2015; Eykholt et al., 2018b; Shafahi et al., 2018; Hosseini & Poovendran, 2018), particularly regarding visual jailbreaks (Carlini et al., 2024; Qi et al., 2023; Li et al., 2024; Feng et al., 2024). Just as visual adversarial attacks have been observed to differ practically and mechanistically from textual attacks (Schaeffer et al., 2024; Wallace et al., 2021), we argue that audio attacks deserve to be studied separately to image attacks. Image and audio input signals fundamentally differ in that audio signals are inherently sequential and require temporal context, unlike static, frame-based image signals, and thus audio requires time-frequency representations like spectrograms whereas images are processed as 2D spatial data. Moreover, human perceptual tolerance for perturbations in audio and image is given through different sets of constraints. Audio is of particular practical interest as vocal interaction with digital assistants is more natural than typed; indeed the nature of chatbot-based communication emulates a transcribed oral conversation. This gives rise to endless meaningful deployments, such as real-time speech analysis for courtrooms, voice biometrics for authentication, audio-based emotion analysis for mental health monitoring or LLM-powered smart home voice assistants (Mahmood et al., 2025; Koffi, 2023).
@@
Contributions. In this paper, we offer first results of an extensive exploration of ALM jailbreaks on the SALMONN-7B language model (Tang et al., 2024). The goal of this research is to deliver a range of empirical results regarding the potential and limitations of jailbreaks in the audio modality, and most importantly, to offer novel insights into the meaning of these in the textual space. We establish an experimental framework to facilitate the study of audio jailbreaks and design a meaningful evaluation dataset for the selected adversarial task. We show that this method permits highly potent jailbreaks which generalize across different content dimensions. We investigate the striking characteristics of the interpretations of these jailbreaks. We discuss the limitations of the jailbreaks and the significance of the selection of the base audio in the optimization process.
@@ 2 Background
Large Language Models and Alignment. LLMs are deep neural networks designed to model the probability distribution of text sequences. One particular training objective is next-token prediction, where, given a sequence of tokens, the model will predict the most likely token to appear next. Broadly, most models have two objectives: the primary training phase focuses on generation of plausible and meaningful text, while the second objective, alignment, aims to make sure that the generated text is ethical and aligns with user-intended goals (Ouyang et al., 2022; Wei et al., 2022). Like other neural-network-based models, these generative systems are susceptible to adversarial attacks, where carefully crafted inputs exploit model vulnerabilities to produce incorrect or unintended outputs (Szegedy et al., 2014; Biggio et al., 2013). An attack that aims to subvert the model’s alignment is called a jailbreak (Wallace et al., 2021; Ebrahimi et al., 2018; Jia & Liang, 2017). Initial jailbreaks were crafted manually and largely found on internet forums (Shen et al., 2024a); nowadays there exists a myriad of algorithmic textual jailbreak methods (Yi et al., 2024). At the time of writing, the state-of-the-art white-box jailbreak method is Greedy Coordinate Gradient (Zou et al., 2023), which iteratively identifies and modifies input tokens to maximize the likelihood of bypassing the model’s alignment constraints by leveraging gradient information. This technique permitted the generation of a ‘universal’ or ‘prompt-agnostic’ jailbreak prefix, that is, a prefix which subverts the model’s safety constraints no matter which harmful prompt it is pre-pended to. The ‘greedy’ aspect refers to the difficulty of discrete optimization - each update has to select the nearest token representation in a continuous search space. It is difficult to make textual jailbreaks stealthy, as the attack text appears clearly unnatural to a reader.
Vision Language Model Jailbreaks. VLMs process both visual and textual information. Typically, this is achieved by encoding images and text into a shared representation space (which is often the representation space of an underlying language model). However, VLMs open up an attack vector on the underlying aligned LM (Carlini et al., 2024). One is able to generate images, e.g. using projected gradient descent, that jailbreak-aligned LMs where purely-textual methods fail to do so. This is a specific example of a broader range of soft-embedding attacks. Qi et al. (2023) generate adversarial visual perturbations in VLMs that constitute a universal image jailbreak that can effectively break alignment on any malicious prompt. The authors find a surprising efficacy even in misalignment categories that the image was not explicitly optimized for. To date, this type of universal jailbreak has not been proven to exist in the audio modality.
Attacks on Automatic Speech Recognition Systems. While audio classification systems were shown to be, unsurprisingly, vulnerable to adversarial attacks (Lan et al., 2024), a more commonly-occurring audio task is Automatic Speech Recognition (ASR). Initial works demonstrated untargeted attacks on ASR that reduced the general transcription quality (Gong & Poellabauer, 2018; Wu & Rajan, 2022), with targeted attacks soon to follow: Carlini & Wagner (2018) show how to craft adversarial perturbations on an arbitrary audio sample to evoke a transcription of choice; Qin et al. (2019) augment this method to introduce properties desirable in a real-world attack scenario such as robustness to degradation and imperceptibility. Nevertheless, ASR systems and ALMs have different architectures and training heuristics. In particular, ASR systems do not combine audio and text, but rather train the model to extract spoken words from the raw signal directly, whereas ALMs perform the embedding projection and cross attention on all audio and text tokens equally for next token prediction. This means that the target task can be arbitrary within the realm of text generation.
@@
Practical Audio Attacks. In this work, we explore audio jailbreaks not only as a theoretical failure mode for ALMs, but also as a practical threat with real-world safety and security implications. To account for this, we incorporate ideas about stealth and psychacoustics in audio signals from Schönherr et al. (2018). There have been many works showing attacks on deployed systems via the audio modality, for example on personal assistants (Ge et al., 2023), spoken assessment (Raina et al., 2020), and speaker verification systems (Kreuk et al., 2018). Interestingly, select works have shown that it is possible to craft image and audio adversarial examples that can easily be reproduced by humans (Khachaturov et al., 2023; Ahmed et al., 2023), which implies that these attacks could be instantiated in a natural environment.
@@
Audio Language Model Jailbreaks. At the time of writing, there are very few works on jailbreaks for ALMs. Some approaches vocalize harmful textual responses in the audio (Yang et al., 2024; Shen et al., 2024b). A recent work proposes Best-of-N-Jailbreaking: a cross-modal per-prompt black-box jailbreak method which works by repeatedly applying random modality-specific augmentations to a harmful request until a harmful response is achieved (Hughes et al., 2024). Kang et al. (2024) use a dual-phase optimization framework, first optimizing discrete latent representations of audio tokens to bypass model safeguards, then refining the corresponding audio waveform while ensuring it remains stealthy and natural through adversarial and retention loss constraints.
Our novel contribution is that we first show it is possible to generate a prompt-agnostic audio jailbreak which transcends the content of the optimization corpus and captures the notion of toxicity as a whole. We explore the effects of different constraints on this optimization process, realistic degradation of the audio, and the resulting jailbreak effectiveness. Secondly, we share insights regarding the meaning and characteristics of the produced jailbreak audios, which has not been explored thus far.
@@ 3 Experimental Design
3.1 Threat Models
Dual Control. The adversary can control the audio and the textual input. In this scenario the adversary is using the audio channel to provoke misbehavior that cannot be elicited solely via the textual channel, i.e. the audio input essentially ‘unlocks’ the unaligned or toxic output. Example: A customer support chatbot for a financial service company integrates speech-to-text and text processing capabilities to assist users with account-related queries. A malicious user enters a question about another user (personal information attack) which the model should, according to it’s alignment, not provide. The textual input alone is declined by the model. Accompanied by a jailbreak-optimized audio however, the user is able to elicit the undesirable response. We also consider a scenario where stealth is required. Example: Consider the same malicious user having to use the system at a public booth. The adversary can avoid suspicious behavior by using an audio that cannot be identified as malicious.
Single Control. Interaction with the ALM is performed exclusively via an audio channel. The textual system prompt is fixed. This would apply to, e.g. call center bots or voice-controlled IoT devices. Example: A customer calls a banking hotline with the system prompt: You are a helpful but harmless and unopinionated assistant to a bank customer.... Using a special jailbreak audio input, the customer is able to subvert the model’s alignment to produce discriminatory content regarding the plausibility of financial loans for certain demographic groups, which the company may liable for. Again, we also consider the case where the audio input has to be crafted to be stealthy. Example: This might occur if the malicious user is trying to evade surveillance/fraud detection.
3.2 Audio-Language Model
We run our experiments on the SALMON-N 7B parameter model developed by Tsinghua University and ByteDance (Tang et al., 2024). SALMON-N consumes audio by extracting both BEATs features (labels such as ‘Snicker’, ‘Drip’ or ‘Human Sounds’) and Whisper features (which are used for speech transcription) from the audio spectogram. These are then combined by dividing the signal into overlapping chunks and fusing them using a Q-former such that these signals are aligned to the language model input space (Kim et al., 2024). The textual input is processed and embedded for the same language model, with the audio and textual tokens concatenated with a delimiter. Thus, the model performs the cross-attention mechanism on these concatenated mixed-modality input tokens. The underlying language model is Vicuna-7Bv1.5, a chatbot trained by fine-tuning LLaMA-7B on user-shared conversations with ChatGPT (Chiang et al., 2023). Vicuna thus mimics the alignment of GPT-4, which is trained using RLHF according to OpenAI’s usage guidelines. Vicuna does not incorporate any further input/output filtering to ensure safety and is vulnerable to many jailbreak generation methods (Chao et al., 2024). We choose to run our experiments on the SALMON-N model due to its open-source nature and its state-of-the-art performance on a range of tasks such as auditory information-based question answering, emotion recognition, and music and audio captioning. Moreover, SALMON-N includes two popular audio feature extractors in a standard configuration with Vicuna. Other ALMs such as Pengi (Deshmukh et al., 2024) or Qwen (Chu et al., 2023) use a very similar architecture. The dual speech and non-speech feature extractor offers high explainability and lends itself well to our meaningfulness exploration.
3.3 Audio Samples
@@ 3.4 Jailbreak Generation
Our method for generating audio jailbreaks was inspired by Qi et al. (2023). Given a base audio $x_{0}$, a target corpus $t=\{t_{0}..t_{n}\}$, and a fully differentiable model $f$, we perform gradient descent on $x_{0}$ to maximize the probability of the output $t$ by minimizing the cross-entropy loss between the predicted distribution and the target outputs:
@@
$\log P_{f}(t_{i}|x)$ is the probability of the model generating the target sentence $t$ with the input audio $x$. During optimization we use an empty textual prompt, a deliberate decision to make the audio jailbreak prompt-agnostic.
As the target corpus $t$, we use a collection of $66$ derogatory sentences directed towards a victim demographic, a victim gender, and the human race in general (Qi et al., 2023). In each epoch, $8$ of these target sentences are optimized on. The fundamental research questions we pose at this stage is can we optimize any base audio such that, when combined with a harmful textual prompt, it reliably circumvents the model’s alignment mechanisms and elicits toxic output?
3.4.1 Stealth
Some of the scenarios described in Section 3.1 require us to augment the optimization formula to generate stealthy audio inputs. This raises the question how does the efficacy of the optimized jailbreak change as we impose stealth constraints? We investigate three approaches to stealth:
  • Epsilon-Constrained (Qi et al., 2023) Here we constrain the absolute change in each audio value, effectively performing bounded gradient descent. In every epoch update, we clip the modified audio such that
    $\forall i:x_{t+1}[i]=\text{clip}(x_{t}[i]-\eta\nabla_{x}\mathcal{L}(x_{t},t),$
    $x_{0}[i]-\epsilon,x_{0}[i]+\epsilon)$
    We take $\epsilon\in\{0.1,0.01,0.001,0.0001\}$.
  • Frequency-Hiding (Schönherr et al., 2018) A normal human hearing range is around 20-20000Hz. However, when adding noise to audio files, it is of course possible to encode information outside of these boundaries. To hide information in specific frequency ranges we use a band-stop filter, removing frequencies between a lower bound $b_{l}$ and an upper bound $b_{u}$:
    $\hat{x}[f]=\begin{cases}x[f],&\text{if }f<b_{l}\text{ or }f>b_{u},\\ 0,&\text{if }b_{l}\leq f\leq b_{u},\end{cases}$
    where $x[f]$ is the frequency component of the audio at frequency $f$, found using a Fourier transform. We experiment with $(b_{l},b_{u})\in\{(1000,8000),(100,10000),(40,20000),(50,15000)\}$.
  • Prepend (Raina et al., 2024) In this scenario, instead of optimizing noise within the audio, we freeze the base audio and optimize a short, unconstrained audio snippet $p$ which is added as a prefix. The loss is calculated on the output resulting from the concatenation of the prefix and the base audio. Given the length $d$ of the prepend snippet in seconds, we now optimize:
    $p^{*}=\arg\min_{p\in[-1,1]^{16000d}}\mathcal{L}([p\|x],t)$
    We randomly initialize the prefix $p$ and experiment with $d\in\{2,1,0.1,0.01\}$.
3.4.2 Audio-Agnostic Jailbreak Noise
Thus far we have discussed techniques to generate adversarial perturbations tailored to a specific input. A natural question to ask is can we make the adversarial noise not only prompt-agnostic but also audio-agnostic? To this end, we optimize our adversarial noise on several audio files simultaneously, to evaluate whether the resulting perturbation retains its effectiveness when applied to entirely new, unseen audio. Using our $n$ base audios $B=\{x^{1}..x^{n}\}$, we optimize a prepend snippet $p$ using the following loss function:
@@ 3.4.2 Audio-Agnostic Jailbreak Noise
[…]
@@ 3.4.3 Robustness
[…]

### v2

@@ Abstract
The rise of multimodal large language models has introduced innovative human-machine interaction paradigms but also significant challenges in machine learning safety. Audio-language Models (ALMs) are especially relevant due to the intuitive nature of spoken communication, yet little is known about their failure modes. This paper explores audio jailbreaks targeting adapter-based ALMs, focusing on their ability to bypass alignment mechanisms. We construct adversarial perturbations that generalize across prompts, tasks, and even base audio samples, demonstrating the first universal jailbreaks in the audio modality, and show that these remain effective in simulated real-world conditions. Beyond demonstrating attack feasibility, we analyze how ALMs interpret these audio adversarial examples and reveal them to encode imperceptible first-person toxic speech – suggesting that the most effective perturbations for eliciting toxic outputs specifically embed linguistic features within the audio signal. These results have important implications for understanding the interactions between different modalities in ALMs, and offer actionable insights for enhancing defenses against adversarial audio attacks.
@@ 1 Introduction
Large Language Models (LLMs) have proven useful beyond a doubt across various domains since their widespread deployment, significantly enhancing productivity in tasks such as natural language processing, code generation, and creative content creation (Brown, 2020; OpenAI, 2024). However, their vast capabilities pose a considerable challenge in balancing usefulness and harmlessness (Bommasani, 2022). One prominent aspect of AI safety is alignment: ensuring that generated content corresponds to the functional objectives and ethical ideals of human users, minimizing risks of harm, bias, or misuse in real-world applications (Weidinger et al., 2021; Russell, 2022). Despite the development of various methods for alignment, such as reinforcement learning from human feedback (Christiano et al., 2023) and rule-based constraints (Mu et al., 2024), LLM alignment has been shown to be inherently brittle and easy to bypass using adversarial prompts, jailbreak techniques, or context manipulation (Perez et al., 2022; Liu et al., 2024; Wei et al., 2023; Xu et al., 2024).
Humans interact primarily through visual and spoken signals, motivating the development of multimodal models that integrate text, images, and audio to better simulate human-like understanding (Baltrušaitis et al., 2019). Among these, Audio Language Models (ALMs) (Chu et al., 2023) take both audio and text as input, introducing a continuous input channel that enables gradient-based adversarial attacks unconstrained by token boundaries (Eykholt et al., 2018; Jia et al., 2022; Carlini & Wagner, 2018). While visual attacks on Vision-Language Models are well-studied (Carlini et al., 2024; Qi et al., 2023; Li et al., 2024; Feng et al., 2024), we argue that audio attacks merit separate investigation due to modality-specific differences (Schaeffer et al., 2024; Wallace et al., 2021). Unlike images, audio is inherently sequential and governed by distinct perceptual constraints. ALMs enable real-world applications such as voice assistants, emotion recognition, and biometric authentication (Mahmood et al., 2025; Koffi, 2023).
@@
Contributions. In this paper, we present novel results of an extensive exploration of ALM jailbreaks on the SALMONN-7B language model (Tang et al., 2024). We deliver a range of empirical results regarding the potential and limitations of jailbreaks in the audio modality, and most importantly, offer novel insights into the interpretation of these in the textual space. We establish an experimental framework to facilitate the study of audio jailbreaks and design a meaningful evaluation dataset for the selected adversarial task. We are able to show highly potent jailbreaks that generalize across different content dimensions and exhibit striking interpretation characteristics.
@@ 2 Background
Large Language Models and Alignment. Large Language Models (LLMs) are trained primarily via next-token prediction to generate coherent text, followed by alignment tuning to ensure outputs align with human intent and ethical guidelines (Ouyang et al., 2022; Wei et al., 2022). Like other neural networks, LLMs are vulnerable to adversarial attacks – inputs designed to elicit unintended behavior (Szegedy et al., 2014; Biggio et al., 2013). Attacks targeting the alignment objective are known as jailbreaks (Wallace et al., 2021; Ebrahimi et al., 2018; Jia & Liang, 2017). Initially handcrafted (Shen et al., 2024a), jailbreaks are now often generated algorithmically (Yi et al., 2024). The current state-of-the-art white-box method, AmpleGCG-Plus (Kumar et al., 2024), trains an LLM on millions of potent Greedy Coordinate Gradient (GCG) (Zou et al., 2023) data points to efficiently produce universal jailbreak prefixes that generalize across harmful prompts.
@@
Practical Audio Attacks. In this work, we explore audio jailbreaks not only as a theoretical failure mode for ALMs, but also as a practical threat with real-world safety and security implications. To account for this, we incorporate ideas about stealth and psychacoustics in audio signals from Schönherr et al. (2018). There have been many works showing attacks on deployed systems via the audio modality, for example on personal assistants (Ge et al., 2023), spoken assessment (Raina et al., 2020), and speaker verification systems (Kreuk et al., 2018).
@@
Audio Language Model Jailbreaks. At the time of writing, there are very few works on jailbreaks for ALMs. Some approaches vocalize harmful textual responses in the audio (Yang et al., 2024; Shen et al., 2024b). A recent work proposes Best-of-N-Jailbreaking: a cross-modal per-prompt black-box jailbreak method which works by repeatedly applying random modality-specific augmentations to a harmful request until a harmful response is achieved (Hughes et al., 2024). Kang et al. (2024) use a dual-phase optimization framework, first optimizing discrete latent representations of audio tokens to bypass model safeguards, then refining the corresponding audio waveform while ensuring it remains stealthy and natural through adversarial and retention loss constraints. In contrast, our work demonstrates that a simple white-box jailbreak method originally conceived for vision-language models can be adapted to generate prompt-agnostic audio jailbreaks that generalize beyond their training corpus and capture broader notions of toxicity. We systematically examine how optimization constraints and real-world audio degradations affect the robustness and effectiveness of these jailbreaks. Additionally, we provide the first analysis of the interpretability and characteristic features of adversarial audio inputs, offering new insights into how models internalize such perturbations.
@@ 3 Experimental Design
Threat Models. We consider two threat models. (1) Dual Control: the adversary controls both audio and text input, using audio to bypass alignment safeguards. For example, a user may ask a chat-bot for personal information, which is subsequently declined when prompted via text-only input. When paired with a jailbreak audio, the model complies. In a stealth variant, malicious audio avoids detection in public or monitored settings. (2) Single Control: Only the audio channel is controlled by the adversary; the system prompt is fixed. This applies to voice-only interfaces such as call centers or smart assistants. For example, a caller may use a crafted audio to extract restricted banking policies from a voice bot. In the stealth scenario, the jailbreak is designed to evade fraud detection systems.
Audio-Language Model. We conduct our experiments on the open-source SALMON-N 7B model (Tang et al., 2024), which integrates audio and text inputs using BEATs (Chen et al., 2022) and Whisper (Radford et al., 2022) features fused via a Q-former. The underlying language model used is Vicuna-7Bv1.5. We select SALMON-N for its popularity, strong performance on audio-based tasks and dual speech and non-speech feature extractors, which provide high explainability. Architectural details are provided in Appendix A.
@@ 3.1 Jailbreak Generation
We adapt the method from Qi et al. (2023) to generate audio-based jailbreaks. Starting from a base audio $x_{0}$ and a target set of toxic sentences $t={t_{0}\dots t_{n}}$, we optimize $x_{0}$ via gradient descent to maximize the likelihood of generating $t$ under a differentiable model $f$:
@@
To ensure prompt-agnosticity, we use empty textual prompts during optimization. The target set $t$ includes 66 toxic sentences targeting individuals and demographic groups (Qi et al., 2023).
Stealth Constraints. In some settings, jailbreak audio must evade human or system detection. We consider three stealth strategies: (1) Epsilon-constrained optimization, which bounds perturbation magnitudes per sample to limit perceptibility (Qi et al., 2023); (2) Frequency-hiding, which suppresses adversarial noise by filtering it into inaudible frequency bands (Schönherr et al., 2018); and (3) Prepend, which optimizes a short audio prefix while keeping the base audio fixed (Raina et al., 2024). Details and formulaic expressions of these constraints are in Appendix D.
3.1.1 Audio-Agnostic Jailbreaks
To generalize across base audios, we jointly optimize a single prepend snippet $p$ over multiple audio samples $B={x^{1}\dots x^{n}}$:
@@ 3.1.1 Audio-Agnostic Jailbreaks
We evaluate transferability by optimizing $p$ on $n-1$ audios and testing it on the n-th held-out sample. This setup yields prompt- and audio-agnostic jailbreaks that remain effective across diverse inputs.
@@ 3.1.2 Robustness
We evaluate jailbreak effectiveness under common degradations that might occur in practice or as naive defenses. Specifically, we test: (1) Over-the-Air Recording using an iPhone 12 at 4 cm in a quiet room; (2) Intermittent Silence Masking via zeroing out short audio segments to simulate dropouts or edits; (3) Gaussian Noise Removal using simple denoising algorithms; and (4) Band-Pass Filtering to remove frequencies outside a target range.